# Analyzing the Effect of Linguistic Similarity on Cross-Lingual Transfer: Tasks and Experimental Setups Matter

### v2

@@ 5.2.1 Should I pick the most similar language according to one similarity measure?
For each target language and each similarity measure, we calculate the difference between the best score obtained by any source language and the score obtained by the most similar language per the measure. [^13] We only consider the source language candidates for which we can compute a similarity score for the target. For size, we select the language with the most training data.
Source = most similar language size pho inv geo syn gb gen lex chr word* POS deV 88 1011 1011 78 57 811 68 POS 2912 1513 1412 1515 1012 1212 0911 1010 1514 1213 LAS 2114 1312 1313 1316 710 1009 810 809 1616 1113 UAS 2712 1613 1512 1615 809 1311 910 1011 1714 1415 topics b. 1712 1714 1314 1512 1413 912 911 1311 405 topics tr. 1310 1311 $1$ 1109 1009 709 708 2013 304 topics m. 1213 1113 1011 910 809 809 809 1210 913
Src = best src lang for the target in another experiment POS deV POS LAS UAS to.-b. to.-t. to.-m. POS deV 304 406 506 610 712 606 POS 609 205 305 911 1012 1614 LAS 709 102 102 1114 1013 1814 UAS 810 305 102 1215 1012 1714 topics b. 1211 1011 1112 1313 307 1915 topics tr. 1009 808 808 909 306 1813 topics m. 506 403 505 505 707 607
Table 3: Mean performance loss if picking a source language solely based on a given metric (top) or based on its transfer performance in another experiment (bottom). Performance loss in percentage points. Subscripts = standard deviations; “word*” = wor, tri, swt; deV = de Vries et al. (2022).
@@
The patterns for the most strongly correlated similarity measures are similar – choosing a source language based on the pertinent similarity measure incurs relatively small losses. E.g., parsing performance is most strongly correlated with syntactic similarity, and picking the syntactically most similar language as the source also results in the lowest performance loss for parsing (Table 3, top). Our findings suggest choosing a syntactically similar language for POS tagging and dependency parsing, and a dataset with high trigram overlap for the $n$-gram-based topic classification experiments.
@@
However, some of the less strong predictors in the global correlations are nearly as good for picking a source language, e.g., genetic and lexical similarity for parsing. We hypothesize that this is due some linguistic measures being more correlated when similarity is high. [^14] E.g., very closely related languages tend to be syntactically similar, even if the overall correlation between genetic and syntactic similarity is only moderately high due the existence of languages that are unrelated but share syntactic similarities or that are (more distantly) related and syntactically dissimilar (Figure 6 in §G). Conversely, character overlap is a worse measure for selecting a source language than the overall correlations would suggest. This is likely due to transfer between languages with low character overlap generally working poorly, while high overlap does not guarantee good transfer (e.g., Figure 7 in Appendix §G).
@@
We additionally simulate a set-up where a researcher has data in multiple source language candidates at their disposal and can afford training and comparing a small selection of models. We select the top three most similar languages according to each measure and take the highest transfer score produced by any of them (Table 12 in §G). The same trends still hold as when picking only one candidate. However, the overall results are much better, and the gaps between the performance losses between the different measures become smaller. Comparing the results of multiple top source language candidates often yields better results than only considering the most similar one.
@@ 5.2.2 Should I choose a source language based on another experiment?
We repeat these analyses but select source languages based on how well they served as source languages for the target language in other experiments (Table 3, bottom; Table 13 in §G).
For the tasks included in our experiments, choosing a source language based on the results for a similar task with similar input representations leads to only small losses: POS tagging results serve as a good predictor for parsing and vice versa, but not for topic classification. Similarly, the results for some topic classification settings are good predictors for each other, but not for POS tagging and parsing. These losses are often even smaller than when selecting source languages based on similarity measures (§5.2.1, Table 3, top).
@@ 6 Conclusion
Despite the fact that our study includes more languages than any prior work, we were able to include only 266 of the world’s $\sim$7000 languages. Moreover, the usual high-resource languages are also over-represented in our work. We take this as an opportunity to encourage data collection in more languages, especially those that are currently extremely low-resource or not available at all.
@@
We investigate one model type per task. Although the choice of our models is well-motivated, we cannot guarantee that our findings hold for other models that could be used for the task. This is an important avenue for investigation in future work.
We do not consider model- or tokenization-specific biases that may have different effects based on the word order White and Cotterell (2021), morphology Park et al. (2021), or writing system and orthography Sproat and Gutkin (2021).
We include $n$-gram-based MLPs in order to be able to have full control over the language model training data. We acknowledge that these models do not have access to word order or syntactic information of the input data, and that they are more affected by exact word choices than a model with (sub)word embeddings would be. We nevertheless include these models because of their computational cheapness, and because there are no (comparable) monolingual embeddings for all $\sim$200 languages in the topic classification data.
@@
The transfer results are overall rather symmetric (i.e., the scores when training on language A and evaluating on language B tend to be similar to when training on B and testing on A; compare the upper right and lower left triangles of the result heatmaps in Figure 2). Nevertheless, these trends are not perfectly symmetrical. Transfer asymmetries have also been observed by other researchers Malkin et al. (2022); Protasov et al. (2024). However, we do not assume that these asymmetries mean that certain languages are intrinsically well-suited as source or target languages, but rather that they reflect peculiarities of the data sets (cf. Bjerva, 2024). We encourage further research on the (a)symmetries of cross-lingual transfer patterns.
Appendix A Languages and Datasets
@@
Table 4: Languages and datasets used in our experiments. Continued in next table.
@@
Tables 4 and 5 show the languages and datasets included in our experiments. Their geographic distribution is pictured in Figure 1.
@@ A.1 Excluded UD treebanks
We exclude treebanks with code-switched language data (qaf_arabizi, qfn_fame, qtd_sagt, qte_tect), without text (ar_nyuad, en_gumreddit, gun_dooley, ja_bccwj, ja_bccwjluw), with glossed language (swl_sslc), and where the division into training and test split changed between releases 2.12 and 2.14 (tr_imst). We also exclude test sets with less than 20 sentences (kfm_aha, nap_rb, nyq_aha, soj_aha) and training sets with less than 100 sentences (bxr_bdt, hsb_ufal, kk_ktb, kmr_mg, lij_glt, olo_kkpp). We additionally exclude other training datasets that were not included in UDPipe 2 (ky_ktmu, de_lit, de_pud). Finally, we exclude the Czech UDPipe-2 models, as they use embeddings from a pretrained Czech model rather than mBERT.
@@ A.2 Excluded SIB-200 languages
We exclude three Arabic dialects whose sentences in SIB-200 are almost identical to the Modern Standard Arabic (arb_Arab) sentences: ars_Arab, acm_Arab, acq_Arab. This is a known issue for FLORES-200, from which SIB-200 is derived.151515github.com/openlanguagedata/flores/issues/8
@@
Appendix B Resources Used
Table 6 lists the versions of the datasets, models, and software we use.
Lang2vec Littell et al. (2017) contains information on syntax, phonology, and phoneme inventories from WALS Dryer and Haspelmath (2013), SSWL Collins and Kayne (2011), Ethnologue Lewis et al. (2015), and PHOIBLE Moran et al. (2014). SIB-200 Adelani et al. (2024) is an annotated subset of FLORES-200 (NLLB Team et al., 2022), which in turn builds on previous versions and extensions of FLORES (Goyal et al., 2022; Guzmán et al., 2019; Doumbouya et al., 2023; AI4Bharat et al., 2023). UDPipe 2 (Straka, 2023) was trained on Universal Dependencies 2.12.161616http://hdl.handle.net/11234/1-5150 Its input representations include the last four layers of base-size uncased mBERT Devlin et al. (2019). This project uses the universal romanizer software ‘uroman’ written by Ulf Hermjakob, USC Information Sciences Institute (2015-2020).
Appendix C Similarity Measures
C.1 Phylogenetic relatedness
For the Glottolog-based phylogenetic distances, we ignore languages associated with the pseudo-families for unclassifiable, unattested, and mixed languages, as well as the “bookkeeping” pseudo-family. For the pseudo-family of constructed languages, we ignore the top-most level (“artificial languages”) and instead retrieve phylogenetic information starting at the second level. We similarly start at the second level for pidgins/creoles.
@@ C.1 Phylogenetic relatedness
C.2 Correlations between similarity measures
@@
Appendix D Hyperparameters and Standard Deviations of Topic Classification Models
@@
Appendix E NLP Task Results
# Datasets # Langs Dataset level Lang level Train/Test Tr/Test Within Across Within  Across POS tagging (accuracy; de Vries et al., 2022) POS 65/105* 65/105 94.1* 4.5 57.4* 22.4 94.1 4.5 57.4 22.4 Grammatical tasks (our experiments) POS 124/268 70/153 96.4 3.1 43.6 20.8 93.4 6.8 43.0 20.2 UAS 88.7 6.4 37.3 19.3 84.8 10.1 36.8 18.8 LAS 84.6 8.6 21.2 17.9 78.4 13.7 20.6 17.0 Topic classification (accuracy; Adelani et al., 2024) MLP 205/205 197/197 62.3 XLM-RB 70.9 XLM-RL 76.1 Topic classification (accuracy; our experiments) base 202/202 194/194 70.3 4.0 20.7 8.8 66.7 14.1 20.1 8.8 trans. 199/199 192/192 69.4 4.1 22.5 8.2 66.7 10.8 22.5 8.2 mb. 202/202 194/194 60.1 18.7 42.6 20.4 58.2 19.9 42.6 20.4
Table 9: Mean scores (in %) of models evaluated within vs. across datasets and within vs. across languages. Standard deviations are formatted as subscripts. *When multiple datasets were available for one language, de Vries et al. (2022) combined them into a single, larger dataset. XLM-RB and L = base- and large-size versions of XLM-R.
Table 9 shows the mean scores within and across languages and datasets for all tasks.
For our topic classification models, the within-dataset performance is comparable to the baselines in the dataset paper (Adelani et al., 2024), but we do not match the performance of their best model.
E.1 Heatmaps of transfer results
@@
Figure 8 shows the POS tagging accuracies for all language pairs. Figures 9 (LAS) and 10 (UAS) present the parsing scores. Figure 11 shows the topic classification accuracies for topics-base, and Figure 12 for topics-translit. Finally, Figure 13 shows the results for the mBERT-based topic classification model.
E.2 Robustness across datasets of the same language
@@
Most treebank pairs of the same language also show very similar transfer patterns. Again, most correlations are close to 1.181818For the training treebanks, we see correlations ($r$) below $0.8$ only for the following pairs: the two German training sets (hdt & gsd: $r=0.75$), the two Persian ones (perft & seraji: $r=0.73$), and the two Korean ones (gsd & kaist: $r=0.72$). For the test datasets this applies to some of the six Latin treebanks (circse & udante: $r=0.79$, circse & ittb: $r=0.79$), several of the eight Turkish datasets (atis & tourism: $r=0.55$, pud & tourism: $r=0.73$, penn & tourism: $r=0.74$, boun & tourism: $r=0.74$, gb & tourism: $r=0.78$, kenet & tourism: $r=0.79$), several of the seven Mandarin Chinese treebanks (hk & patentchar: $r=0.77$, beginner & patentchar: $r=0.78$, patentchar & pud: $r=0.79$, cfl & patentchar: $r=0.79$, gsd & patentchar: $r=0.79$), two of the four Old East Slavic treebanks (birchbark & rnc: $r=0.75$), the two Akkadian treebanks (pisandub & riao: $r=0.75$), and both Sanskrit treebanks (ufal & vedic: $r=0.57$).
@@
E.3 Effect of writing system
@@
Nonetheless, writing systems still play a role, especially for the $n$-gram-based models. Although topics-base and topics-translit achieve the same within-language accuracy (Tables 2/9), topics-translit performs slightly better cross-lingually. Its within-dataset performance is slightly lower than for topics-base (69.4% vs. 70.3%), likely due to some language-specific information being lost when diacritics are removed. This would be compensated in the within-language performance by improved transfer between datasets of languages that originally had different scripts.
E.3.1 Effects of transliteration
@@
Appendix F Correlations
F.1 Correlations between task results
@@
Table 10: Correlations between task results (Pearson’s r). Correlations between results of the same task types (POS/parsing vs. topic classification) are on a dataset level; correlations between different task types are on a language level. All p-values are below 0.0001.
@@
F.2 Mixed effects models
@@
Table 11: Linear mixed effects model results for each experiment. P-values are based on model comparison: *** = < 0.001, ** = < 0.01, * = < 0.05, . = < 0.1. Entries with p-values of >=0.05 are in grey. The last two rows show the number of training and test languages included in the analysis (i.e., the language pairs for which no fixed effects had missing information). “Corr” indicates pairs of strongly correlated fixed effects (see text, §F.2). “Same_script” is True iff the training and test datasets use the same writing system; “mbert_test” is True iff the test language is one of mBERT’s pretraining languages.
@@
F.3 Correlations between NLP results and similarity measures
Tables 14 and 15 show the correlations between POS/parsing scores and the similarity measures for each test language. Tables 16, 17, and 18 show the same for the topic classification experiments. Because these tables take up a lot of space, they are placed at the very end of the appendix.
Appendix G Implications for Choosing Source Languages
size pho inv geo syn gb gen lex chr word* Top-1 source candidate POS deV 88 1011 1011 78 57 811 68 POS 2912 1513 1412 1515 1012 1212 0911 1010 1514 1213 LAS 2114 1312 1313 1316 710 1009 810 809 1616 1113 UAS 2712 1613 1512 1615 809 1311 910 1011 1714 1415 topics b. 1712 1714 1314 1512 1413 912 911 1311 405 topics tr. 1310 1311 $1$ 1109 1009 709 708 2013 304 topics m. 1213 1113 1011 910 809 809 809 1210 913 Top-3 source candidates POS deV 34 45 45 25 24 58 24 POS 2513 708 708 507 304 406 507 406 708 507 LAS 1814 809 709 508 304 406 406 305 811 609 UAS 2413 808 708 608 305 506 507 507 910 709 topics b. 1010 911 509 607 609 507 306 809 204 topics tr. 808 709 408 506 506 305 305 1411 103 topics m. 607 505 405 404 405 404 405 504 611
Table 12: Mean performance loss (in percentage points) if picking source languages based solely on a given metric. Standard deviations are written in subscripts. “Word*” refers to word overlap for the UD tasks, trigram overlap for the n-gram topic classification models, and subword token overlap for the mBERT-based MLP. Top: Loss when only considering the single most similar language (or the language with the most training data). Bottom: Loss when comparing the top-3 most similar languages (/languages with the most training data).
POS deV POS LAS UAS to. b. to. tr. to. m. Top-1 source candidate POS deV 304 406 506 610 712 606 POS 609 205 305 911 1012 1614 LAS 709 102 102 1114 1013 1814 UAS 810 305 102 1215 1012 1714 topics b. 1211 1011 1112 1313 307 1915 topics tr. 1009 808 808 909 306 1813 topics m. 506 403 505 505 707 607 Top-3 source candidates POS deV 103 204 304 305 203 404 POS 306 102 103 406 406 707 LAS 305 001 508 508 910 UAS 406 103 001 508 507 809 topics b. 608 507 508 508 103 1211 topics tr. 506 506 405 506 205 1210 topics m. 202 202 202 202 303 304
Table 13: Mean performance loss (in percentage points) if picking source languages based solely on performance on the task in the column. Standard deviations are written in subscripts. Top: Loss when only considering the single best-performing language on the other task. Bottom: Loss when comparing the top-3 best-performing languages on the other task.
This section provides additional context for §5.2.1. Table 12 shows the mean performance loss when choosing source languages based on their similarity to the target language (or based on their training data size). Table 12 shows the loss when choosing the source languages that were the best sources in another cross-lingual transfer task.
Figure 6 illustrates the relationship between genetic and syntactic similarity – phylogenetically unrelated and distantly related languages run the gamut of syntactic similarity scores, but all closely related languages are syntactically rather similar. We hypothesize that this is why of some of the variables with relatively weak or moderate overall correlations are still good predictors when choosing a single source language (or a small set of source languages). For instance, while phylogenetic similarity shows much weaker correlations with parsing scores than syntactic similarities overall, it is a good proxy when picking a single source, since closely related languages also tend to have similar syntax.
Figure 7 shows the relationship between character overlap between training and test sets and the classification scores of the topics-base model. Character overlap is a worse metric for selecting a source language than the overall correlations would suggest. This is likely due to transfer between languages with low character overlap generally working poorly, while high overlap does not guarantee good transfer.
Figure 6: Phylogenetic relatedness scores compared to syntactic relatedness scores.
Figure 7: Character overlap between topic classification training and test sets compared to topic classification accuracy (topics-base).

### v3

@@ 5.2.1 Should I pick the most similar language according to one similarity measure?
For each target language and each similarity measure, we calculate the difference between the best score obtained by any source language and the score obtained by the most similar language per the measure. We only consider the source language candidates for which we can compute a similarity score for the target. For size, we select the language with the most training data. The patterns for the most strongly correlated similarity measures are similar – choosing a source language based on the pertinent similarity measure incurs relatively small losses. E.g., parsing performance is most strongly correlated with syntactic similarity, and picking the syntactically most similar language as the source also results in the lowest performance loss for parsing (Table 3, top). Our findings suggest choosing a syntactically similar language for POS tagging and dependency parsing, and a dataset with high trigram overlap for the $n$-gram-based topic classification experiments.
Figure 4: Left: Relationship between phylogenetic and syntactic similarity – unrelated or distantly related languages can be syntactically similar or dissimilar, but all closely related languages are syntactically similar. Right: Relationship between character overlap (between training and test sets) and the classification scores of the topics-base model – transfer between languages with low character overlap works poorly, but high overlap does not guarantee good transfer.
@@
However, some of the less strong predictors in the global correlations are nearly as good for picking a source language, e.g., genetic and lexical similarity for parsing. We hypothesize that this is due some linguistic measures being more correlated when similarity is high. [^11] E.g., very closely related languages tend to be syntactically similar, even if the overall correlation between genetic and syntactic similarity is only moderately high due the existence of languages that are unrelated but share syntactic similarities or that are (more distantly) related and syntactically dissimilar (Figure 4, left). Conversely, character overlap is a worse measure for selecting a source language than the overall correlations would suggest. This is likely due to transfer between languages with low character overlap generally working poorly, while high overlap does not guarantee good transfer (e.g., Figure 4, right).
@@
We additionally simulate a setup where a researcher has data in multiple source language candidates at their disposal and can afford training and comparing a small selection of models. We select the top three most similar languages according to each measure and take the highest transfer score produced by any of them (Table 3, bottom). The same trends still hold as when picking only one candidate. However, the overall results are much better, and the gaps between the performance losses between the different measures become smaller. Comparing the results of multiple top source language candidates often yields better results than only considering the most similar one.
POS deV POS LAS UAS to.-b. to.-t. to.-m. Top-1 source candidate (= best src in another experiment) POS deV 304 406 506 610 712 606 POS 609 205 305 911 1012 1614 LAS 709 102 102 1114 1013 1814 UAS 810 305 102 1215 1012 1714 top.-b. 1211 1011 1112 1313 307 1915 top.-tr. 1009 808 808 909 306 1813 top.-m. 506 403 505 505 707 607 Top-3 source candidates POS deV 103 204 304 305 203 404 POS 306 102 103 406 406 707 LAS 305 001 508 508 910 UAS 406 103 001 508 507 809 top.-b. 608 507 508 508 103 1211 top.-tr. 506 506 405 506 205 1210 top.-m. 202 202 202 202 303 304
Table 4: Mean performance loss if picking source languages based solely on performance on the task in the column. Performance loss in percentage points. Subscripts = standard deviations; “word*” = wor, tri, swt; deV = de Vries et al. (2022).
@@ 5.2.2 Should I choose a source language based on another experiment?
We repeat these analyses but select source languages based on how well they served as source languages for the target language in other experiments Table 4).
For the tasks included in our experiments, choosing a source language based on the results for a similar task with similar input representations leads to only small losses: POS tagging results serve as a good predictor for parsing and vice versa, but not for topic classification. Similarly, the results for some topic classification settings are good predictors for each other, but not for POS tagging and parsing. These losses are often even smaller than when selecting source languages based on similarity measures (cf. Tables 3 and 4).
@@ 6 Conclusion
Acknowledgements
We would like to thank Katherine Metcalf, Maureen de Seyssel, Sinead Williamson, Skyler Seto, Stéphane Aroca-Ouellette, and the anonymous reviewers for their helpful comments, discussions, and feedback.
@@ Acknowledgements
Despite the fact that our study includes more languages than any prior work, we were able to include only 263 of the world’s $\sim$7000 languages. Moreover, the usual high-resource languages are also over-represented in our work. We encourage data collection in more languages, especially those that are currently extremely low-resource or not available at all.
@@
We investigate one model type per task. Our model choices are motivated by practical considerations. We use UDPipe in part because of its state-of-the-art status for many languages and its wide coverage of languages, making it a likely out-of-the-box tool to be used for (morpho)syntactic annotations. An important motivation for us to choose MLPs for our topic classification experiments is the lightweight architecture that allows us to quickly train and compare many models, and we show that the performance is not far off from larger, more resource-intensive models (Table 2). We include $n$-gram-based MLPs in order to be able to have full control over the language model training data. Despite these models not having access to word order or syntactic information of the input data (and being more affected by exact word choices than a model with (sub)word embeddings would be), we include them because of their computational cheapness, and because there are no (comparable) monolingual embeddings for all $\sim$200 languages in the topic classification data.
We cannot guarantee that our findings hold for other models that could be used for the task, which is an avenue for investigation in future work. We also do not consider model- or tokenization-specific biases that may have different effects based on the word order White and Cotterell (2021), morphology Park et al. (2021), or writing system and orthography Sproat and Gutkin (2021).
@@
The transfer results are overall rather symmetric (i.e., the scores when training on language A and evaluating on language B tend to be similar to when training on B and testing on A; compare the upper right and lower left triangles of the result heatmaps in Figure 2). However, these trends are not perfectly symmetrical. Transfer asymmetries have also been observed by other researchers Malkin et al. (2022); Protasov et al. (2024). However, we do not assume that these asymmetries mean that certain languages are intrinsically well-suited as source or target languages, but rather that they reflect peculiarities of the data sets (cf. Bjerva, 2024). We encourage further research on the (a)symmetries of cross-lingual transfer patterns.
Appendix A Languages and Resources
Table 5 lists the versions of the datasets, models, and software we use. Lang2vec Littell et al. (2017) contains information on syntax, phonology, and phoneme inventories from WALS Dryer and Haspelmath (2013), SSWL Collins and Kayne (2011), Ethnologue Lewis et al. (2015), and PHOIBLE Moran et al. (2014). SIB-200 Adelani et al. (2024) is an annotated subset of FLORES-200 (NLLB Team et al., 2022), which in turn builds on previous versions and extensions of FLORES (Goyal et al., 2022; Guzmán et al., 2019; Doumbouya et al., 2023; AI4Bharat et al., 2023). UDPipe 2 (Straka, 2023) was trained on Universal Dependencies 2.12.121212http://hdl.handle.net/11234/1-5150 Its input representations include the last four layers of base-size uncased mBERT Devlin et al. (2019). This project uses the universal romanizer software ‘uroman’ written by Ulf Hermjakob, USC Information Sciences Institute (2015-2020).
@@ Appendix A Languages and Resources
Table 6: Languages and datasets used in our experiments. Continued and explained in next table.
@@
Tables 6 and 7 show the languages and datasets included in our experiments. Their geographic distribution is pictured in Figure 1.
@@ A.1 Excluded UD treebanks
We exclude treebanks with code-switched language data (qaf_arabizi, qfn_fame, qtd_sagt, qte_tect), without text (ar_nyuad, en_gumreddit, gun_dooley, ja_bccwj, ja_bccwjluw), with glossed language (swl_sslc), and where the division into training and test split changed between releases 2.12 and 2.14 (tr_imst). We also exclude test sets with fewer than 20 sentences (kfm_aha, nap_rb, nyq_aha, soj_aha) and training sets with less than 100 sentences (bxr_bdt, hsb_ufal, kk_ktb, kmr_mg, lij_glt, olo_kkpp). We additionally exclude other training datasets that were not included in UDPipe 2 (ky_ktmu, de_lit, de_pud). Finally, we exclude the Czech UDPipe-2 models, as they use embeddings from a pretrained Czech model rather than mBERT.
@@ A.2 Excluded SIB-200 languages
We exclude three Arabic dialects whose sentences in SIB-200 are almost identical to the Modern Standard Arabic (arb_Arab) sentences: ars_Arab, acm_Arab, acq_Arab. This is a known issue for FLORES-200, from which SIB-200 is derived.131313github.com/openlanguagedata/flores/issues/8
@@
Appendix B Correlations between similarity measures
@@
Appendix C Hyperparameters and Standard Deviations of Topic Classification Models
@@
Appendix D NLP Task Results
D.1 Random baselines
For POS tagging and topic classification, we consider random baselines that randomly predict one of the 17 POS tags (or one of the seven topics) and are thus correct 5.9 % (or 14.3 %) of the time.
For each setup, all within-language performances (the diagonals in Figure 2) are above this threshold. Thus, all models learned something about the task, regardless of the training language and could be expected to be able to transfer some of that knowledge to a well-suited test language.
For some of the train–test language combinations, we see performances that are worse than random chance. This is the case for 3.1 % of the POS tagging results, 26.6 % of topics-base’s results, 14.5 % of topics-translit’s results, and 6.8 % of the results by topics-mbert. These worse-than-random transfer results are also meaningful: a POS tagger (or parser) trained on one language might learn grammatical patterns that run counter to how another language works, or a topic classification model might be mislead by superficial string overlaps between datasets that are false friends.
D.2 Heatmaps of transfer results
@@
Figure 7 shows the POS tagging accuracies for all language pairs. Figures 8 (LAS) and 9 (UAS) present the parsing scores. Figure 10 shows the topic classification accuracies for topics-base, and Figure 11 for topics-translit. Finally, Figure 12 shows the results for the mBERT-based topic classification model.
D.3 Robustness across datasets of the same language
@@
Most treebank pairs of the same language also show very similar transfer patterns. Again, most correlations are close to 1.151515For the training treebanks, we see correlations ($r$) below $0.8$ only for the following pairs: the two German training sets (hdt & gsd: $r=0.75$), the two Persian ones (perft & seraji: $r=0.73$), and the two Korean ones (gsd & kaist: $r=0.72$). For the test datasets this applies to some of the six Latin treebanks (circse & udante: $r=0.79$, circse & ittb: $r=0.79$), several of the eight Turkish datasets (atis & tourism: $r=0.55$, pud & tourism: $r=0.73$, penn & tourism: $r=0.74$, boun & tourism: $r=0.74$, gb & tourism: $r=0.78$, kenet & tourism: $r=0.79$), several of the seven Mandarin Chinese treebanks (hk & patentchar: $r=0.77$, beginner & patentchar: $r=0.78$, patentchar & pud: $r=0.79$, cfl & patentchar: $r=0.79$, gsd & patentchar: $r=0.79$), two of the four Old East Slavic treebanks (birchbark & rnc: $r=0.75$), the two Akkadian treebanks (pisandub & riao: $r=0.75$), and both Sanskrit treebanks (ufal & vedic: $r=0.57$).
@@
D.4 Effect of writing system
@@
Nonetheless, writing systems still play a role, especially for the $n$-gram-based models. Although topics-base and topics-translit achieve the same within-language accuracy (Table 2), topics-translit performs slightly better cross-lingually. Its within-dataset performance is slightly lower than for topics-base (69.4% vs. 70.3%), likely due to some language-specific information being lost when diacritics are removed. This would be compensated in the within-language performance by improved transfer between datasets of languages that originally had different scripts.
D.4.1 Effects of transliteration
@@
Appendix E Correlations
E.1 Correlations between task results
@@
Table 10: Correlations between task results (Pearson’s r, all p-values are below 0.0001). Where possible, correlations are on a dataset level, otherwise on a language level.
@@
E.2 Mixed effects models
@@
Table 11: Linear mixed effects model results for each experiment. P-values are based on model comparison: *** = < 0.001, ** = < 0.01, * = < 0.05, . = < 0.1. Entries with p-values of >=0.05 are in grey. The last two rows show the number of training and test languages included in the analysis (i.e., the language pairs for which no fixed effects had missing information). “Corr” indicates pairs of strongly correlated fixed effects (see text, §E.2). “Same_script” is True iff the training and test datasets use the same writing system; “mbert_test” is True iff the test language is one of mBERT’s pretraining languages.
@@
E.3 Correlations between NLP results and similarity measures
Tables 12 and 13 show the correlations between POS/parsing scores and the similarity measures for each test language. Tables 14, 15, and 16 show the same for the topic classification experiments. Because these tables take up a lot of space, they are placed at the very end of the appendix.